3) Find the indicated probability. A die with 12 sides is rolled. What is the probability of rolling a number less than 11?
Answer #1

Solution:-

The probability is 5/6 = 0.8333

Explanation:- There are 10 number below 11 on the die,so the probability

P(A) = number of favorable condition/total number of condition
= 10/12
= 5/6
= 0.8333